The price elasticity for rice is estimated to be -0.4 per pound and income elasticity is 0.8. At a price of $0.40 per pound and a per capita income of $20,000, the demand for rice is 50 Million tons per year.

a) is rice an inferior good, a necessity or a luxury? Explain
b)if per capita income increases to $20,500, what will be the quantity demanded of rice?
c) if the price of rice increases to $0.41 per pound and income per capita remains at $20,000, what will be quantity demanded?
1
Expert's answer
2015-02-27T09:05:45-0500
The price elasticity -0.4, income elasticity is 0.8. P = $0.40 per pound, I = $20,000, Qd = 50 Million tons per year.
a) As income elasticity of demand for rice is 0.8 (E < 1), rice is an inferior good.
b) If per capita income increases to $20,500, the change in quantity demanded of rice will be 0.8*(20,500 - 20,000)/20,000 = 0.02 or +2%.
So, the new quantity demanded is Qd 50*1.02 = 51 million tons.
c) If the price of rice increases to $0.41 per pound and income per capita remains at $20,000, the change in quantity demanded is -0.4*(0.41 - 0.40)/0.40 = -1, so the new quantity demanded is Qd = 50 - 1 = 49 million tons.